(2, -8) and (x, 2); slope: -10/3
Answer by Alan3354(69443)

(2, -8) and (x, 2); slope: -10/3
----
Slope m = diffy/diffx
m = (-8-2)/(2 - x) = -10/3
-10/(2-x) = -10/3
-30 = -10(2-x) = -20 + 10x
10x = -10
x = -1